Vector A is in the direction 34.0* clockwise from the
Y axis. The x-component of A is Ax= -16.0m. (a) What is the Y-component Of A? (b) What is the magnitude of A?
1
Expert's answer
2019-10-10T09:54:46-0400

The components of vector "\bold a" are as follows: "a \sin \theta = - a_x", "a \cos \theta = a_y", where "a = |\bold a|", "\theta = 34^{\circ}", "a_x = 16".

From the first equation, "a = \frac{-a_x}{\sin \theta}", and plugging it into the second equation, obtain "a_y = -a_x \ctg \theta \approx 23.72".

The magnitude of "\bold a" is "a = \sqrt{a_x^2 + a_y^2} \approx 28.61".